# An operator learning perspective on parameter-to-observable maps

### v1

@@ Abstract.
Computationally efficient surrogates for parametrized physical models play a crucial role in science and engineering. Operator learning provides data-driven surrogates that map between function spaces. However, instead of full-field measurements, often the available data are only finite-dimensional parametrizations of model inputs or finite observables of model outputs. Building off of Fourier Neural Operators, this paper introduces the Fourier Neural Mappings (FNMs) framework that is able to accommodate such finite-dimensional inputs and outputs. The paper develops universal approximation theorems for the method. Moreover, in many applications the underlying parameter-to-observable (PtO) map is defined implicitly through an infinite-dimensional operator, such as the solution operator of a partial differential equation. A natural question is whether it is more data-efficient to learn the PtO map end-to-end or first learn the solution operator and subsequently compute the observable from the full-field solution. A theoretical analysis of Bayesian nonparametric regression of linear functionals, which is of independent interest, suggests that the end-to-end approach can actually have worse sample complexity. Extending beyond the theory, numerical results for the FNM approximation of three nonlinear PtO maps demonstrate the benefits of the operator learning perspective that this paper adopts.
@@
The first author is supported by the high-performance computing platform of Peking University. The second author acknowledges support from the National Science Foundation Graduate Research Fellowship Program under award number DGE-1745301 and from the Amazon/Caltech AI4Science Fellowship, and partial support from the Air Force Office of Scientific Research under MURI award number FA9550-20-1-0358 (Machine Learning and Physics-Based Modeling and Simulation). The third author is supported by the Department of Energy Computational Science Graduate Fellowship under award number DE-SC00211. The second and third authors are also grateful for partial support from the Department of Defense Vannevar Bush Faculty Fellowship held by Andrew M. Stuart under Office of Naval Research award number N00014-22-1-2790.
$*{}^{*}$orresponding author: Nicholas H. Nelsen (nnelsen@caltech.edu)
@@
${}^{1}$eijing International Center for Mathematical Research, Peking University, China
${}^{2}$omputing and Mathematical Sciences, California Institute of Technology, USA
@@ 1. Introduction
Operator learning has emerged as a methodology that enables the machine learning of maps between spaces of functions. Many surrogate modeling tasks in areas such as uncertainty quantification, inverse problems, and design optimization involve a map between function spaces, such as the solution operator of a partial differential equation (PDE). However, the primary quantities of interest (QoI) in these tasks are usually just a finite number of design parameters or output observables. This may be because full-field data, such as initial conditions, boundary conditions, and solutions of PDEs, are not accessible from measurements or are too expensive to acquire. The prevailing approach then involves emulating the parameter-to-observable (PtO) map instead of the underlying solution map between function spaces. Yet, it is natural to wonder if the success of operator learning in the function-to-function setting can be brought to bear in this more realistic setting where inputs or outputs may necessarily be finite-dimensional. To this end, the present paper introduces Fourier Neural Mappings (FNMs) as a way to extend operator learning architectures such as the Fourier Neural Operator (FNO) to finite-dimensional input and output spaces in a manner compatible with the underlying operator between infinite-dimensional spaces. The admissible types of FNM models considered in this work are visualized in Figure 1.
Nevertheless, it is possible to accommodate finite-dimensional inputs or outputs through other means. For instance, one could lift a finite-dimensional input to a function by expanding in predetermined basis functions, apply traditional operator learning architectures to the full-field function space data, and then directly compute a known finite-dimensional QoI from the output function. In contrast, the end-to-end FNM approach in this work is fully data-driven and operates directly on finite-dimensional data without the need for pre- and postprocessing. A natural question is whether one of these two approaches achieves better accuracy than the other when the goal is to predict certain QoIs. In the present paper, we address this important question both from a theoretical and a numerical perspective. Indeed, it has been empirically observed in various nonlinear problems ranging from electronic structure calculations [73] to metamaterial design [5] that data-driven methods that predict the full-field response of a system are superior to end-to-end approaches for the same downstream tasks or QoIs. The analysis in this paper takes the first steps toward a rigorous theoretical understanding of these empirical findings.
Throughout the paper, we refer to learning a function-valued map as full-field learning. Given such a learned map, various known QoIs may be directly computed from the output of the map. On the other hand, we refer to the direct estimation of the map from an input to the observed QoI as end-to-end learning. This terminology distinguishes between output spaces. When either the input or output is finite- and the other is infinite-dimensional, we label it as “vector-to-function” (V2F) or “function-to-vector” (F2V), respectively, to avoid ambiguity. The abbreviations V2V and F2F for “vector-to-vector” and “function-to-function” are analogous.
@@ 1.1. Contributions
    We introduce FNMs as a function space architecture that is able to accommodate finite-dimensional inputs, outputs, or both.
@@
    We perform numerical experiments with FNMs in three different examples—an advection–diffusion equation, flow over an airfoil, and an elliptic homogenization problem—that show empirical evidence that the theoretical linear intuition from Contribution (C3) remains valid for nonlinear maps.
@@ 1.2. Related work
Several works have established neural operators as a viable tool for scientific machine learning. The general neural operator formalism is described in [31] and contains several subclasses including DeepONet [48], graph neural operator [40, 42], and FNO [41]. These architectures allow for function data evaluated at different grid points or resolutions to be used with the same model. In particular, the FNO is primarily parametrized in Fourier space. It exploits the fact that the Fourier basis spans $L^{2}$ and uses the efficient Fast Fourier Transform (FFT) algorithm for computations. The idea of parametrizing operators in Fourier space is explored in earlier works as well [54, 56]. The FNO has been shown to be applicable both to domains other than the torus and to nonuniform meshes [39, 45]. These neural operators have been used in various areas of application, including climate modeling [32], fracture mechanics [22], and catheter design [81]. In several of these applications, neural operators have been implemented with finite-dimensional inputs or outputs by using constant functions as substitutes for finite vectors, which is theoretically justified by statements of universal approximation [7], or by using other hand-designed maps. However, learning a constant function as a representation for a constant is unnatural and computationally wasteful; it is desirable to substitute a more suitable architecture. The present paper develops FNMs that extend neural operators to this important setting while retaining desirable universal approximation properties.
@@
The theory of scientific machine learning falls broadly into three tiers. In the first tier, universal approximation results [13, 14, 25, 76] use classical approximation theory to guarantee that the architecture is capable of representing maps from within a class of interest to any desired accuracy. Some of the proofs of these results contain constructive arguments, but the corresponding architectures are usually not as empirically effective as those that solely come with existence results. Examples of constructive arguments for operator approximation are contained in [23], which constructs ReLU neural networks, and [8], which uses randomized numerical linear algebra to sketch Green’s functions for linear elliptic PDEs. Each of these works also comes equipped with convergence rates with respect to model size and data size, respectively; these rates form the second tier of operator learning theory. Many papers in this tier prove bounds on the required model size, i.e., parameter complexity [17, 24, 30, 33, 35, 37, 47, 67]. Some are able to obtain sample complexity bounds, although most results are restricted to linear or kernelized settings [11, 16, 27, 36, 43, 53, 66, 71]. The third tier of theory describes the likelihood of actually obtaining an accurate approximation through optimization. While some results along these lines exist for linear models, linear maps, and constructive operators [36, 38], they are absent for the class of neural operators optimized through variants of stochastic gradient descent (SGD). This is the class that has proven empirically most effective in applications thus far and is the class used in this work.
@@
To provide theoretical intuition for nonlinear settings, this paper establishes convergence rates in the tractable setting of learning a linear functional—the PtO map—from noisy data. The setting of functional linear regression has a long history in statistics [9, 12, 63, 78]. A zoo of different estimators exist; e.g., those based on reproducing kernel Hilbert space (RKHS) methods, principal component analysis, and wavelets. We target the frequentist convergence properties of a linear Gaussian posterior estimator that arises from reinterpreting the regression task as a Bayesian inverse problem. The closest work to ours is [44]. There, the authors derive posterior contraction rates for Bayesian functional linear regression with Gaussian priors. However, their main results are only sharp if the prior is correctly specified to match the regularity of the true linear functional. Our work goes beyond this by proving sharp high-probability error bounds for out-of-distribution prediction error under very general smoothness misspecification. However, to do so we require that the prior and data covariance operators commute, which is a limitation of our theory. Another relevant work is [10], in which the authors make minimal assumptions on the prior and data covariances but still make the well-specified assumption that the truth belongs to the RKHS of the prior. They also require a data-dependent scaling of the prior. Similarly, [57] obtains convergence guarantees for Bayesian nonparametric regression of functions with Gaussian process priors in the misspecified setting. However, their work only considers a squared exponential covariance structure and requires a careful rescaling of the prior to deal with the smoothness misspecification. Our work holds for the larger class of misspecified Matérn-like Gaussian priors and delivers rates without rescaling the prior distribution.
@@
Beyond linear functionals, recent work proposes and analyzes a kernelized deep learning method for nonlinear functionals [68]. The idea of such neural functionals, a subclass of the FNMs proposed in this work, is not new. One appearance is in the context of a function space discriminator for generative adversarial networks [60]. However, that work uses only a single bounded linear functional that is appended to the output of a FNO and is parametrized by a standard neural network. This is a special case of our FNMs for function-to-vector maps. Another paper that shares similar ideas to ours is [80]. There, the authors also formulate a V2V neural network approach that maps through a latent 1D function space. However, their encoder and decoder maps are prescribed by hand-picked basis functions, while for FNMs the encoder and decoder maps are learned from data.
@@
In this paper, three potential applications are highlighted. The first application is an advection-diffusion model where the input is a velocity field and the output is the state at a fixed future time. This problem is considered a benchmark for scientific machine learning [72]. Some theoretical approximation rates for it have been developed for DeepONet in the F2F setting [17]. The second application centers on the compressible flow over an airfoil, i.e., a wing cross section. This experiment is explored for FNO in [39] and used as a shape optimization example in the F2F setting for DeepONet in [69] and for reduced basis networks in [55]. Several other related works devise V2V-based neural network approaches and novel training strategies for this aerodynamics problem [49, 50, 51, 52]. The third application involves learning the effective homogenized elasticity coefficient for a multiscale elliptic PDE. This example is explored in detail for FNO in [7] and for other constitutive laws in [46]. For the Darcy flow—or scalar coefficient—setting of this equation, other work has adopted the F2F setting to efficiently compute QoIs [77]. For each of these applications, we compare the generalization error performance of all four F2F, F2V, V2F, and V2V variants of FNMs as well as standard fully-connected neural networks.
@@ 1.3. Notation
The set of continuous linear operators from a Banach space $\mathcal{U}$ to a Banach space $\mathcal{V}$ is denoted by $\mathcal{L}(\mathcal{U};\mathcal{V})$, and if $\mathcal{U}=\mathcal{V}$, then we write $\mathcal{L}(\mathcal{V})$. For a separable Hilbert space $(H,\langle\cdot,\cdot\rangle,\lVert{\,\cdot\,}\rVert)$, the outer product operator $a\otimes b\in\mathcal{L}(H)$ is defined by $(a\otimes b)c\coloneqq\langle b,c\rangle a$ for any elements $a$, $b$, and $c$ of $H$. Let $(\Omega,\mathcal{F},\operatorname{\mathbb{P}})$ be a probability space that is sufficiently rich to support all random variables that appear in this paper. All expectations are interpreted as Bochner integrals. For a probability measure $\Pi$ supported on $H$ with two finite moments, we denote by $\operatorname{Cov}(\Pi)=\operatorname{\mathbb{E}}^{u\sim\Pi}[(u-\operatorname{\mathbb{E}}u)\otimes(u-\operatorname{\mathbb{E}}u)]\in\mathcal{L}(H)$ its covariance operator. Independent and identically distributed (iid) random variables $X_{1},X_{2},\ldots,X_{n}$ from $\Pi$ are denoted by $\{X_{i}\}_{i=1}^{n}\sim\Pi^{\otimes n}$. For two random variables $X$ and $Z$, the conditional expectation notation $\operatorname{\mathbb{E}}^{Z\,|\,X}[{\,\cdot\,}]$ denotes integration with respect to the law of $Z\,|\,X$. For $N\in\mathbb{N}$, we write $[N]\coloneqq\{1,2,\ldots,N\}$. For two nonnegative real sequences $\{a_{n}\}_{n\in\mathbb{N}}$ and $\{b_{n}\}_{n\in\mathbb{N}}$, we write $a_{n}\lesssim b_{n}$ if there exists $c>0$ such that $a_{n}\leq cb_{n}$ for all $n\in\mathbb{N}$ and $a_{n}\simeq b_{n}$ if both $a_{n}\lesssim b_{n}$ and $b_{n}\lesssim a_{n}$. To denote asymptotic equivalence, we write $a_{n}\asymp b_{n}$ as $n\to\infty$ if there exist $c\geq 1$ and $n_{0}\in\mathbb{N}$ such that $c^{-1}b_{n}\leq a_{n}\leq cb_{n}$ for all $n\geq n_{0}$. The Sobolev-like sequence Hilbert spaces $\mathcal{H}^{s}=\mathcal{H}^{s}(\mathbb{N},\mathbb{R})$ are defined for $s\in\mathbb{R}$ by $\mathcal{H}^{s}\coloneqq\{v\colon\mathbb{N}\to\mathbb{R}\,|\,\sum_{j=1}^{\infty}j^{2s}\lvert v_{j}\rvert^{2}<\infty\}$. We write $\mathbb{T}^{d}$ for the $d$-dimensional unit torus $[0,1]_{\mathrm{per}}^{d}$.

### v2

@@ Abstract.
Computationally efficient surrogates for parametrized physical models play a crucial role in science and engineering. Operator learning provides data-driven surrogates that map between function spaces. However, instead of full-field measurements, often the available data are only finite-dimensional parametrizations of model inputs or finite observables of model outputs. Building on Fourier Neural Operators, this paper introduces the Fourier Neural Mappings (FNMs) framework that is able to accommodate such finite-dimensional vector inputs or outputs. The paper develops universal approximation theorems for the method. Moreover, in many applications the underlying parameter-to-observable (PtO) map is defined implicitly through an infinite-dimensional operator, such as the solution operator of a partial differential equation. A natural question is whether it is more data-efficient to learn the PtO map end-to-end or first learn the solution operator and subsequently compute the observable from the full-field solution. A theoretical analysis of Bayesian nonparametric regression of linear functionals, which is of independent interest, suggests that the end-to-end approach can actually have worse sample complexity. Extending beyond the theory, numerical results for the FNM approximation of three nonlinear PtO maps demonstrate the benefits of the operator learning perspective that this paper adopts.
@@
Corresponding author: Nicholas H. Nelsen (nnelsen@caltech.edu)
@@
1Beijing International Center for Mathematical Research, Peking University, China
2Computing and Mathematical Sciences, California Institute of Technology, USA
@@ 1. Introduction
Operator learning has emerged as a methodology that enables the machine learning of maps between spaces of functions. Many surrogate modeling tasks in areas such as uncertainty quantification, inverse problems, and design optimization involve a map between function spaces, such as the solution operator of a partial differential equation (PDE). However, the primary quantities of interest (QoI) in these tasks are usually just a finite number of design parameters or output observables. This may be because full-field data, such as initial conditions, boundary conditions, and solutions of PDEs, are not accessible from measurements or are too expensive to acquire. The prevailing approach then involves emulating the parameter-to-observable (PtO) map instead of the underlying solution map between function spaces. Yet, it is natural to wonder if the success of operator learning in the function-to-function setting can be brought to bear in this more realistic setting where inputs or outputs may necessarily be finite-dimensional vectors. To this end, the present paper introduces Fourier Neural Mappings (FNMs) as a way to extend operator learning architectures such as the Fourier Neural Operator (FNO) to finite-dimensional input and output spaces in a manner compatible with the underlying operator between infinite-dimensional spaces. The admissible types of FNM models considered in this work are visualized in Figure 1.
@@
Nevertheless, it is possible to accommodate finite-dimensional inputs or outputs through other means. For instance, one could lift a finite-dimensional input vector to a function by expanding in predetermined basis functions, apply traditional operator learning architectures to the full-field function space data, and then directly compute a known finite-dimensional QoI from the output function. In contrast, the end-to-end FNM approach in this work is fully data-driven and operates directly on finite-dimensional vector data without the need for pre- and postprocessing. A natural question is whether one of these two approaches achieves better accuracy than the other when the goal is to predict certain QoIs. In the present paper, we address this important question both from a theoretical and a numerical perspective. Indeed, it has been empirically observed in various nonlinear problems ranging from electronic structure calculations [84] to metamaterial design [6] that data-driven methods that predict the full-field response of a system are superior to end-to-end approaches for the same downstream tasks or QoIs. The analysis in this paper takes the first steps toward a rigorous theoretical understanding of these empirical findings.
Throughout the paper, we refer to learning a function-valued map as full-field learning. Given such a learned map, various known QoIs may be directly computed from the output of the map. On the other hand, we refer to the direct estimation of the map from an input to the observed QoI as end-to-end learning. This terminology distinguishes between output spaces. When either the input or output is a finite vector and the other is infinite-dimensional, we label the learning approach as “vector-to-function” (V2F) or “function-to-vector” (F2V), respectively, to avoid ambiguity. The abbreviations V2V and F2F for “vector-to-vector” and “function-to-function” are analogous.
@@ 1.1. Contributions
    We introduce FNMs as a function space architecture that is able to accommodate finite-dimensional vector inputs, outputs, or both.
@@
    We perform numerical experiments with FNMs in three different examples—an advection–diffusion equation, flow over an airfoil, and an elliptic homogenization problem—that show empirical evidence that the theoretical linear insight from Contribution (C3) remains valid for nonlinear maps.
@@ 1.2. Related work
Several works have established neural operators as a viable tool for scientific machine learning. The general neural operator formalism is described in [37] and contains several subclasses including DeepONet [56], graph neural operator [47, 49], and FNO [48]. These architectures allow for function data evaluated at different grid points or resolutions to all be used with the same model. In particular, the FNO is primarily parametrized in Fourier space. It exploits the fact that the Fourier basis spans $L^{2}$ and uses the efficient Fast Fourier Transform (FFT) algorithm for computations. The idea of parametrizing operators in Fourier space is explored in earlier works as well [63, 67]. The FNO has been shown to be applicable both to domains other than the torus and to nonuniform meshes [46, 52]. These neural operators have been used in various areas of application, including climate modeling [38], fracture mechanics [24], and catheter design [92]. In several of these applications, neural operators have been implemented with finite-dimensional vector inputs or outputs by using constant functions as substitutes for finite vectors, which is theoretically justified by statements of universal approximation [8], or by using other hand-designed maps. However, learning a constant function as a representation for a constant is unnatural and computationally wasteful; it is desirable to substitute a more suitable architecture. The present paper develops FNMs that extend neural operators to this important setting while retaining desirable universal approximation properties.
@@
The theory of scientific machine learning falls broadly into three tiers. In the first tier, universal approximation results [14, 15, 29, 20] use classical approximation theory to guarantee that the architecture is capable of representing maps from within a class of interest to any desired accuracy. Some of the proofs of these results contain constructive arguments, but the corresponding architectures are usually not as empirically effective as those that solely come with existence results. Examples of constructive arguments for operator approximation are contained in [27], which constructs ReLU neural networks, and [9], which uses randomized numerical linear algebra to sketch Green’s functions for linear elliptic PDEs. Each of these works also comes equipped with convergence rates with respect to model size and data size, respectively; these rates form the second tier of operator learning theory. Many papers in this tier prove bounds on the required model size, i.e., parameter complexity [18, 28, 35, 39, 41, 43, 55, 78]. Some are able to obtain sample complexity bounds, although most results are restricted to linear or kernelized settings [12, 17, 31, 42, 50, 54, 61, 77, 82]. The third tier of theory describes the likelihood of actually obtaining an accurate approximation through optimization. While some results along these lines exist for linear models, linear maps, and constructive operators [42, 45, 64], they are absent for the class of neural operators optimized through variants of stochastic gradient descent (SGD). This is the class that has proven empirically most effective in applications thus far and is the class used in this work.
@@
To provide theoretical intuition for nonlinear settings, this paper establishes convergence rates in the tractable setting of learning a linear functional—the PtO map—from noisy data. The setting of functional linear regression has a long history in statistics [10, 13, 74, 89]. A zoo of different estimators exist, e.g., those based on reproducing kernel Hilbert space (RKHS) methods, principal component analysis, and wavelets. We target the frequentist convergence properties of a linear Gaussian posterior estimator that arises from reinterpreting the regression task as a Bayesian inverse problem. The closest work to ours is [51]. There, the authors derive posterior contraction rates for Bayesian functional linear regression with Gaussian priors. However, their main results are only sharp if the prior is correctly specified to match the regularity of the true linear functional. Our work goes beyond this by proving sharp high-probability error bounds for out-of-distribution prediction error under very general smoothness misspecification. However, to do so we require that the prior and data covariance operators commute, which is a limitation of our theory. Another relevant work is [11], in which the authors make minimal assumptions on the prior and data covariances but still make the well-specified assumption that the truth belongs to the RKHS of the prior. They also require a data-dependent scaling of the prior. Similarly, [68] obtains convergence guarantees for Bayesian nonparametric regression of functions with Gaussian process priors in the misspecified setting. However, their work only considers a squared exponential covariance structure and requires a careful rescaling of the prior to deal with the smoothness misspecification. Our work holds for the larger class of misspecified Matérn-like Gaussian priors and delivers rates without rescaling the prior distribution.
@@
Beyond linear functionals, recent work proposes and analyzes a kernelized deep learning method for nonlinear functionals [79]. The idea of such neural functionals, a subclass of the FNMs proposed in this work, is not new. One appearance is in the context of a function space discriminator for generative adversarial networks [71]. However, that work uses only a single bounded linear functional that is appended to the output of a FNO and is parametrized by a standard neural network. This is a special case of our FNMs for F2V maps. Another paper that shares similar ideas to ours is [91]. There, the authors also formulate a V2V neural network approach that maps through a latent 1D function space. However, their encoder and decoder maps are prescribed by hand-picked basis functions, while for FNMs the encoder and decoder maps are learned from data.
@@
In this paper, three potential applications are highlighted. The first application is an advection–diffusion model where the input is a velocity field and the output is the state at a fixed future time. This problem is considered a benchmark for scientific machine learning [83]. Some theoretical approximation rates for it have been developed for DeepONet in the F2F setting [18]. The second application centers on the compressible flow over an airfoil, i.e., a wing cross section. This experiment is explored for FNO in [46] and used as a shape optimization example in the F2F setting for DeepONet in [80] and for reduced basis networks in [66]. Several other related works devise V2V-based neural network approaches and novel training strategies for this aerodynamics problem [57, 58, 59, 60]. The third application involves learning the homogenized elasticity coefficient for a multiscale elliptic PDE. This example is explored in detail for FNO in [8] and for other constitutive laws in [53]. For the Darcy flow—or scalar coefficient—setting of this equation, other work adopts the F2F setting to efficiently compute QoIs [88]. For each of these applications, we compare the generalization error performance of all four F2F, F2V, V2F, and V2V variants of FNMs as well as standard fully-connected neural networks.
@@ 1.3. Notation
The set of continuous linear operators from a Banach space $\mathcal{U}$ to a Banach space $\mathcal{V}$ is denoted by $\mathcal{L}(\mathcal{U};\mathcal{V})$, and if $\mathcal{U}=\mathcal{V}$, then we write $\mathcal{L}(\mathcal{V})$. For a separable Hilbert space $(H,\langle\cdot,\cdot\rangle,\lVert{\,\cdot\,}\rVert)$, the outer product operator $a\otimes b\in\mathcal{L}(H)$ is defined by $(a\otimes b)c\coloneqq\langle b,c\rangle a$ for any elements $a$, $b$, and $c$ of $H$. Let $(\Omega,\mathcal{F},\operatorname{\mathbb{P}})$ be a probability space that is sufficiently rich to support all random variables that appear in this paper. All expectations are interpreted as Bochner integrals. For a probability measure $\Pi$ supported on $H$ with two finite moments, we denote by $\operatorname{Cov}(\Pi)=\operatorname{\mathbb{E}}^{u\sim\Pi}[(u-\operatorname{\mathbb{E}}u)\otimes(u-\operatorname{\mathbb{E}}u)]\in\mathcal{L}(H)$ its covariance operator. Independent and identically distributed (i.i.d.) random variables $X_{1},X_{2},\ldots,X_{n}$ from $\Pi$ are denoted by $\{X_{i}\}_{i=1}^{n}\sim\Pi^{\otimes n}$. For two random variables $X$ and $Z$, the conditional expectation notation $\operatorname{\mathbb{E}}^{Z\,|\,X}[{\,\cdot\,}]$ denotes integration with respect to the law of $Z\,|\,X$. For $N\in\mathbb{N}$, we write $[N]\coloneqq\{1,2,\ldots,N\}$. For two nonnegative real sequences $\{a_{n}\}_{n\in\mathbb{N}}$ and $\{b_{n}\}_{n\in\mathbb{N}}$, we write $a_{n}\lesssim b_{n}$ if there exists $c>0$ such that $a_{n}\leq cb_{n}$ for all $n\in\mathbb{N}$ and $a_{n}\simeq b_{n}$ if both $a_{n}\lesssim b_{n}$ and $b_{n}\lesssim a_{n}$. To denote asymptotic equivalence, we write $a_{n}\asymp b_{n}$ as $n\to\infty$ if there exist $c\geq 1$ and $n_{0}\in\mathbb{N}$ such that $c^{-1}b_{n}\leq a_{n}\leq cb_{n}$ for all $n\geq n_{0}$. The Sobolev-like sequence Hilbert spaces $\mathcal{H}^{s}=\mathcal{H}^{s}(\mathbb{N},\mathbb{R})$ are defined for $s\in\mathbb{R}$ by $\mathcal{H}^{s}\coloneqq\{v\colon\mathbb{N}\to\mathbb{R}\,|\,\sum_{j=1}^{\infty}j^{2s}\lvert v_{j}\rvert^{2}<\infty\}$. We write $\mathbb{T}^{d}$ for the $d$-dimensional unit torus $[0,1]_{\mathrm{per}}^{d}$.